# Search or Accelerate: Confidence-Switched Position Beam Search for Diffusion Language Models

### v1

@@ Abstract
Diffusion Language Models (DLMs) generate text by iteratively denoising a masked sequence, repeatedly deciding which positions to commit at each step. Standard decoding follows a greedy rule, unmask the most confident positions, yet this local choice can lock the model into a suboptimal unmasking order, especially on reasoning-heavy prompts. We present Search Or AcceleRate (SOAR), a training-free decoding algorithm that adapts its behavior to the model’s uncertainty. When confidence is low, SOAR briefly widens the search over alternative unmasking decisions to avoid premature commitments; when confidence is high, it collapses the search and decodes many positions in parallel to reduce the number of denoising iterations. Across mathematical reasoning and code generation benchmarks (GSM8K, MBPP, HumanEval) on Dream-7B and LLaDA-8B, SOAR improves generation quality while maintaining competitive inference speed, offering a practical way to balance quality and efficiency in DLM decoding.
@@ 2.2 Optimized Inference Methods for DLLMs
The non-autoregressive, iterative nature of DLM decoding presents unique inference challenges, as the random access pattern across denoising steps invalidates the traditional KV-cache and necessitates pre-defining a maximum sequence length, often leading to computational redundancy. Consequently, research has focused on specialized acceleration techniques. One primary direction involves designing specialized KV-cache mechanisms that either leverage the high similarity of hidden states across steps for approximate caching or restructure decoding into a block-autoregressive manner to enable state reuse from previous context (Wu et al., 2025a; Nguyen-Tri et al., 2025; Huang et al., 2025a; Liu et al., 2025; Ma et al., 2025). Another critical direction aims to reduce the total number of steps through parallel and adaptive-length decoding. This includes confidence-based strategies that dynamically determine how many tokens to decode in parallel, as well as early-commit methods which halt decoding once the model’s predictions stabilize, significantly speeding up inference (Wu et al., 2025a; Yang et al., 2025b; Gao et al., 2025; Li et al., 2025b, a). In addition, some works addresses the inherent inability of standard DLMs to revise committed tokens, exploring solutions ranging from training-free correctors to modifications of the diffusion process itself (Huang et al., 2025b; Mounier and Idehpour, 2025; Zhu et al., 2025b). A concurrent work, Order-Token Search (Anonymous, 2025), proposes to simultaneously explore the decoding space across both the position and token dimensions to achieve better results, but at the expense of inference speed. Our proposed method, SOAR, dynamically switches between search and parallel decoding modes based on the model’s confidence during token decoding. This enables consistent improvements across multiple benchmarks without sacrificing decoding speed.
@@ 3.1 Preliminaries of DLM Inference
where $\mathcal{M}(\cdot,\rho_{t})$ masks a proportion $\rho_{t}$ of tokens, and $\delta(\cdot)$ is the Dirac delta function. The reverse process iteratively reconstructs the original text:
$p_{\theta}(x_{t-1}|x_{t})=\prod_{i=1}^{L}p_{\theta}(x_{t-1,i}|x_{t})$ (2)
@@ 3.2.1 Candidate Generation
Using the average instead of the sum in the scoring function aims to prevent sequences with more remaining masked tokens from being unfairly penalized.
@@ 4.1 Experiment Setup
We conduct experiments on two representative diffusion language models: LLaDA-8B (Nie et al., 2025) and Dream-7B (Ye et al., 2025), both using the Base version. To ensure reproducibility, all experiments are conducted with a temperature of 0, using softmax probabilities as confidence scores. Unless otherwise specified, the threshold $\tau=0.95$ for LLaDA-8B and $\tau=0.90$ for Dream-7B and the maximum beam size K is set to 2. The maximum sequence length is set to $256$ and $512$, respectively, to validate the robuestness of the results.
@@
We adhere to standard few-shot evaluation protocols for each benchmark: 0-shot for HumanEval, 3-shot for MBPP and 4-shot for GSM8K. Accuracy is used as the evaluation metric for mathematical reasoning, while pass@1 is adopted for code generation benchmarks. All experiments are conducted based on open-source lm-evaluation-harness with a single NVIDIA A100 80GB GPU. The total inference time of the benchmark is used to caculate the SpeedUp ratio compared with standard greedy decoding.
Figure 4: The threshhold study on Dream-7B: (a) Cumulative distribution of token confidence scores on GSM8K; (b) Trade-off between accuracy and SpeedUp under varying confidence thresholds on MBPP; (c) Trade-off between accuracy and SpeedUp under varying confidence thresholds on GSM8K.
@@ 4.5 Analysis of the Robustness of SOAR
Table 3 compares the performance of standard greedy decoding and SOAR across different initial sequence lengths. The results demonstrate that SOAR, as a training-free method, consistently maintains or improves performance when integrated with DreamOn’s variable-length decoding mechanism. This experiment confirms SOAR ’s flexibility and robustness, highlighting its potential to be effectively combined with other carefully designed decoding strategies without requiring architectural modifications or additional training.
@@ 5 Conclusion
We introduce Search Or AcceleRate (SOAR), an adaptive inference framework that dynamically balances generation quality and speed for Diffusion Language Models. By allowing the model to widen its search when uncertain and accelerate when confident, SOAR consistently improves generation quality across coding and mathematical reasoning tasks while maintaining competitive inference speed. Extensive experiments confirm its robustness to different unmask metrics, sequence lengths, and base models. As a training-free method, SOAR offers a practical and effective solution to the quality-speed trade-off in DLM inference.
@@ Appendix A Visualization of SOAR
Increasingly Parallel Decoding. We analyzed the usage patterns of Beam Search and Parallel Decoding modes across decoding steps on the Dream-7B-Base model using GSM8K. Consistent with the processing in Subsection 4.4, for all sample decoding sequences, we retain only the valid portion — i.e., the tokens preceding the keyword “answer”. Since different samples have varying effective decoding lengths $S_{\text{sample}}$, we record the ratio of each token’s decoding step to its effective sequence length as a measure of the Decoding Process(%). Figure 7 illustrates how the two decoding modes evolve as decoding progresses. In the early stages, SOAR allocates a larger proportion of steps (55%) to PBS to explore more promising sequences. As decoding advances, the model eventually favors parallel decoding in 85% of the cases. This trend demonstrates the model’s transition from initial uncertainty toward greater confidence. Throughout the inference process, SOAR dynamically switches between the two decoding modes based on confidence scores in an adaptive manner.

### v2

@@ Abstract
Diffusion Language Models (DLMs) generate text by iteratively denoising a masked sequence, repeatedly deciding which positions to commit at each step. Standard decoding follows a greedy rule, unmask the most confident positions, yet this local choice can lock the model into a suboptimal unmasking order, especially on reasoning-heavy prompts. We present Search Or AcceleRate (SOAR), a training-free decoding algorithm that adapts its behavior to the model’s uncertainty. When confidence is low, SOAR briefly widens the search over alternative unmasking decisions to avoid premature commitments; when confidence is high, it collapses the search and decodes many positions in parallel to reduce the number of denoising iterations. Across mathematical reasoning and code generation benchmarks (GSM8K, MBPP, HumanEval) on Dream-7B and LLaDA-8B, SOAR improves generation quality while maintaining competitive inference speed, offering a practical way to balance quality and efficiency in DLM decoding. Our Code is available at https://github.com/duterscmy/SOAR
@@ 2.2 Optimized Inference Methods for DLLMs
The non-autoregressive, iterative nature of DLM decoding presents unique inference challenges, as the random access pattern across denoising steps invalidates the traditional KV-cache and necessitates pre-defining a maximum sequence length, often leading to computational redundancy. Thus, research has focused on specialized acceleration techniques.
One primary direction involves designing specialized KV-cache mechanisms that either leverage the high similarity of hidden states across steps for approximate caching or restructure decoding into a block-autoregressive manner to enable state reuse from previous context (Wu et al., 2025a; Nguyen-Tri et al., 2025; Huang et al., 2025a; Liu et al., 2025; Ma et al., 2025). Another critical direction aims to reduce the total number of steps through parallel and adaptive-length decoding. This includes confidence-based strategies that dynamically determine how many tokens to decode in parallel, as well as early-commit methods which halt decoding once the model’s predictions stabilize, significantly speeding up inference (Wu et al., 2025a; Yang et al., 2025b; Gao et al., 2025; Li et al., 2025b, a). In addition, some works address the inherent inability of standard DLMs to revise committed tokens, exploring solutions ranging from training-free correctors to modifications of the diffusion process itself (Huang et al., 2025b; Mounier and Idehpour, 2025; Zhu et al., 2025b). A concurrent work, Order-Token Search (Anonymous, 2025), proposes to simultaneously explore the decoding space across both the position and token dimensions to achieve better results, but at the expense of inference speed. Our proposed method, SOAR, dynamically switches between search and parallel decoding modes based on the model’s confidence during token decoding. This enables consistent improvements across multiple benchmarks without sacrificing decoding speed.
@@ 3.1 Preliminaries of DLM Inference
where $\mathcal{M}(\cdot,\rho_{t})$ masks a proportion $\rho_{t}$ of tokens, and $\delta(\cdot)$ is the Dirac delta function. The reverse process iteratively reconstructs the original text by modeling the reverse transition as conditionally independent across positions in $x_{t-1}$ given the corrupted sequence $x_{t}$:
$p_{\theta}(x_{t-1}|x_{t})=\prod_{i=1}^{L}p_{\theta}(x_{t-1,i}|x_{t}),$ (2)
@@ 3.2.1 Candidate Generation
Using the average instead of the sum in the scoring function ensures that scores remain comparable across sequences with different numbers of remaining masked tokens.
@@ 4.1 Experiment Setup
We conduct experiments on two representative diffusion language models: LLaDA-8B (Nie et al., 2025) and Dream-7B (Ye et al., 2025), both using the Base version. Unless otherwise specified, we use the softmax probabilities as confidence scores. The confidence threshold $\tau$ is set to $0.95$ for LLaDA-8B and $0.90$ for Dream-7B, and the maximum beam size $K$ is set to $2$. The maximum sequence length is set to $256$ and $512$, respectively, to validate the robustness of the results.
@@
We adhere to standard few-shot evaluation protocols for each benchmark: 0-shot for HumanEval, 3-shot for MBPP and 4-shot for GSM8K. Accuracy is used as the evaluation metric for mathematical reasoning, while pass@1 is adopted for code generation benchmarks. All experiments are conducted based on open-source lm-evaluation-harness (Gao et al., 2024) with a single NVIDIA A100 80GB GPU. The total inference time of the benchmark is used to calculate the SpeedUp ratio compared with standard greedy decoding.
Figure 4: The threshold study on Dream-7B: (a) Cumulative distribution of token confidence scores on GSM8K; (b) Trade-off between accuracy and SpeedUp under varying confidence thresholds on MBPP; (c) Trade-off between accuracy and SpeedUp under varying confidence thresholds on GSM8K.
@@ 4.5 Analysis of the Robustness of SOAR
Table 3 compares the performance of standard greedy decoding and SOAR across different initial sequence lengths. The results demonstrate that SOAR, as a training-free method, consistently maintains or improves performance when integrated with DreamOn’s variable-length decoding mechanism. This confirms SOAR ’s flexibility and robustness, highlighting its potential to be effectively combined with other carefully designed decoding strategies without requiring architectural modifications or additional training.
@@ 5 Conclusion
We introduce Search Or AcceleRate (SOAR), an adaptive inference framework that dynamically balances generation quality and speed for DLMs. By allowing the model to widen its search when uncertain and accelerate when confident, SOAR consistently improves generation quality across coding and mathematical reasoning tasks while maintaining competitive inference speed. Ablation studies further validate the robustness of SOAR under varying confidence thresholds, beam sizes, and unmask metrics. Moreover, SOAR generalizes effectively across different base models and sequence lengths, and integrates seamlessly with variable-length decoding strategies like DREAMON. As a training-free method, SOAR offers a practical and effective solution to the quality-speed trade-off in DLM inference.
@@ Appendix A Visualization of SOAR
Increasingly Parallel Decoding. We analyzed the usage patterns of Beam Search and Parallel Decoding modes across decoding steps on the Dream-7B-Base model using GSM8K. Consistent with the processing in Subsection 4.4, for all sample decoding sequences, we retain only the valid portion, i.e., the tokens preceding the keyword “answer”. Since different samples have varying effective decoding lengths $S_{\text{sample}}$, we record the ratio of each token’s decoding step to its effective sequence length as a measure of the Decoding Process (%). Figure 7 illustrates how the two decoding modes evolve as decoding progresses. In the early stages, SOAR allocates a larger proportion of steps (55%) to PBS to explore more promising sequences. As decoding advances, the model eventually favors parallel decoding in 85% of the cases. This trend demonstrates the model’s transition from initial uncertainty toward greater confidence. Such behavior is expected, as the progressive reduction in remaining masked tokens increasingly constrains the conditional distribution over subsequent decoding steps, resulting in a more concentrated distribution and higher model confidence. Throughout the inference process, SOAR dynamically switches between the two decoding modes based on confidence scores in an adaptive manner.
@@
Appendix B Pseudocode of SOAR
Algorithm 1 presents the complete decoding procedure of SOAR, which dynamically switches between parallel decoding and position beam search based on token-level confidence.
Algorithm 1 Search Or AcceleRate (SOAR)
0: Prompt $x_{\text{prompt}}$, DLM $f_{\theta}$, max length $L$, confidence threshold $\tau$, max beam width $B_{\text{max}}$
0: Generated sequence $x_{0}$
1: Initialize $x_{T}$ with $x_{\text{prompt}}$ followed by $L$ [MASK] tokens
2: Initialize beam $\mathcal{B}_{T}=\{(x_{T},0)\}$
3: Set candidate size $N_{T}=1$ {Initial candidate size}
4:for $t=T,T-1,\dots,1$ do
5:  $\mathcal{C}_{t}\leftarrow\emptyset$ {Candidate set for step tt}
6:  for each $(x_{t}^{(j)},s_{t}^{(j)})\in\mathcal{B}_{t}$ do
7:   Compute token distributions $\mathbf{p}_{t}^{(j)}=f_{\theta}(x_{t}^{(j)})$
8:   Compute confidence scores $c_{t,i}^{(j)}=\max_{v}\mathbf{p}_{t,i}^{(j)}[v]$ for all positions $i$
9:   Identify masked positions $M_{t}^{(j)}=\{i:x_{t,i}^{(j)}=\text{[MASK]}\}$
10:   Identify high-confidence positions $H_{t}^{(j)}=\{i\in M_{t}^{(j)}:c_{t,i}^{(j)}>\tau\}$
11:   if $H_{t}^{(j)}\neq\emptyset$ then
12:    {— Parallel Decoding Mode —}
13:    $\mathcal{I}\leftarrow H_{t}^{(j)}$ {Unmask all high-confidence positions}
14:    $x_{t-1}\leftarrow\text{Unmask}(x_{t}^{(j)},\mathcal{I},\mathbf{p}_{t}^{(j)})$
15:    $\Delta s\leftarrow\frac{1}{|\mathcal{I}|}\sum_{i\in\mathcal{I}}c_{t,i}^{(j)}$
16:    Add $(x_{t-1},s_{t}^{(j)}+\Delta s,\textsc{Parallel})$ to $\mathcal{C}_{t}$
17:   else
18:    {— Position Beam Search Mode —}
19:    Sort masked positions by confidence: $i_{1},i_{2},\dots$ where $c_{t,i_{1}}^{(j)}\geq c_{t,i_{2}}^{(j)}\geq\cdots$
20:    $k\leftarrow\min(B_{\text{max}},|M_{t}^{(j)}|)$
21:    for $\ell=1,\dots,k$ do
22:     $x_{t-1}^{(\ell)}\leftarrow\text{Unmask}(x_{t}^{(j)},\{i_{\ell}\},\mathbf{p}_{t}^{(j)})$
23:     Add $(x_{t-1}^{(\ell)},s_{t}^{(j)}+c_{t,i_{\ell}}^{(j)},\textsc{BeamSearch})$ to $\mathcal{C}_{t}$
24:    end for
25:   end if
26:  end for
27:  {— Candidate Pruning and Beam Update —}
28:  Sort $\mathcal{C}_{t}$ by average cumulative score in descending order
29:  Let $(x^{*},s^{*},\text{mode}^{*})$ be the top candidate in $\mathcal{C}_{t}$
30:  if $\text{mode}^{*}=\textsc{Parallel}$ then
31:   {High confidence: focus on the best path}
32:   Set candidate size $N_{t}\leftarrow 1$
33:  else
34:   {Uncertainty detected: maintain multiple hypotheses}
35:   Set candidate size $N_{t}\leftarrow\min(B_{\text{max}},|\mathcal{C}_{t}|)$
36:  end if
37:  $\mathcal{B}_{t-1}\leftarrow$ top $N_{t}$ candidates from $\mathcal{C}_{t}$
38:end for
39:return $x_{0}$ from the best candidate in $\mathcal{B}_{0}$